OpenStudy (anonymous):

Select one of the factors of 3x2 + 4x - 4. (x + 4) (3x + 2) (3x - 2) (x - 2)

hartnn (hartnn):

can you find 2 numbers whose sum is +4 and product is 3(-4) = -12 ?

OpenStudy (anonymous):

wait what? lol i remember something about that but can you explain as to what exactly i have to do?

hartnn (hartnn):

sure :) to factor \(ax^2+bx+c\) we split the middle term : bx by finding 2 numbers with sum as +b and product as ac

hartnn (hartnn):

so for \(3x^2 +4x-4\) we need to split 4x find 2 numbers with sum as +4 and peoduct as -12

hartnn (hartnn):

*product

OpenStudy (anonymous):

I see why you would try to find 4 but why -12?

OpenStudy (anonymous):

oh wait is it becuase you have to multiply 3 & 4

hartnn (hartnn):

yes ac = 3(-4) =-12 so that we would be able to factor out the quadratic into 2 factors properly

OpenStudy (anonymous):

oh okay that makes sense.

hartnn (hartnn):

so... can you find 2 numbers whose sum is +4 and product is 3(-4) = -12 ?

hartnn (hartnn):

hint : one number is negative, other is positive

OpenStudy (anonymous):

6\[6-2=4 and 6\times-2\]

hartnn (hartnn):

good! so we split 4x as \(\large 4x =6x -2x\)

hartnn (hartnn):

\(\Large 3x^2 +6x-2x-4\) what can you factor out from 1st 2 terms \(3x^2 +6x\) ??

OpenStudy (anonymous):

sorry i had to answer the phone

hartnn (hartnn):

\(\Large 3x^2 = 3x (x) \\ \Large 6x = 3x(2)\) no problem!

OpenStudy (anonymous):

this is where i get stuck. so would you do this?

OpenStudy (anonymous):

\[3x(x+2)?\]

hartnn (hartnn):

yes!

hartnn (hartnn):

so, now take last 2 terms -2x-4 what can you factor out ?

OpenStudy (anonymous):

-2(x-2)?

hartnn (hartnn):

-4 is actually -2(+2) not -2(-2)

OpenStudy (anonymous):

oops sorry

hartnn (hartnn):

so \(\Large -2x-4 = -2 (x+2)\) makes sense ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

we have now \(\large 3x\color{blue}{(x+2)} -2 \color {blue}{(x+2)}\) what can be factored out ??

OpenStudy (anonymous):

(x+2)(3x-2)?

hartnn (hartnn):

\(\Huge \checkmark \)

OpenStudy (anonymous):

awesome thank you!!!

hartnn (hartnn):

welcome ^_^